OpenStudy (error1603):

Solve the system of equations - 3 + 5x = y 2x + 5y = 12

OpenStudy (error1603):

@Whitemonsterbunny17 @mathmale

OpenStudy (error1603):

@.Sam.

OpenStudy (error1603):

@ganeshie8

OpenStudy (error1603):

@King.Void. @Elsa213

OpenStudy (wanell):

x,y} = {7,6} System of Linear Equations entered : [1] 2x + y = 20 [2] 6x - 5y = 12 Graphic Representation of the Equations : y + 2x = 20 -5y + 6x = 12

OpenStudy (wanell):

/ Solve equation [1] for the variable y [1] y = -2x + 20 // Plug this in for variable y in equation [2] [2] 6x - 5•(-2x+20) = 12 [2] 16x = 112 // Solve equation [2] for the variable x [2] 16x = 112 [2] x = 7 // By now we know this much : x = 7 y = -2x+20 // Use the x value to solve for y y = -2(7)+20 = 6 Solution : {x,y} = {7,6}

OpenStudy (error1603):

Thanks